# XEQ Scale for Evaluating XAI Experience Quality Grounded in Psychometric Theory

### v2

@@ Abstract
Explainable Artificial Intelligence (XAI) aims to improve the transparency of autonomous decision-making through explanations. Recent literature has emphasised users’ need for holistic “multi-shot” explanations and the ability to personalise their engagement with XAI systems. We refer to this user-centred interaction as an XAI Experience. Despite advances in creating XAI experiences, evaluating them in a user-centered manner has remained challenging. To address this, we introduce the XAI Experience Quality (XEQ) Scale (pronounced "Seek" Scale), for evaluating the user-centered quality of XAI experiences. Furthermore, XEQ quantifies the quality of experiences across four evaluation dimensions: learning, utility, fulfilment and engagement. These contributions extend the state-of-the-art of XAI evaluation, moving beyond the one-dimensional metrics frequently developed to assess single-shot explanations. In this paper, we present the XEQ scale development and validation process, including content validation with XAI experts as well as discriminant and construct validation through a large-scale pilot study. Out pilot study results offer strong evidence that establishes the XEQ Scale as a comprehensive framework for evaluating user-centred XAI experiences.
@@ 1 Introduction
Explainable Artificial Intelligence (XAI) describes a range of techniques to elucidate autonomous decision-making and the data that informed that AI system [19, 12, 2]. Each technique typically provides explanations that focus on a specific aspect of the system and its decisions. Accordingly, the utility of employing multiple techniques for a holistic explanation of a system becomes increasingly clear [2, 27]. The collection of explanations, provided by different techniques and describing different components of the system, forms what we describe as “multi-shot” explanations. Previous work has demonstrated the effectiveness of delivering multi-shot explanations using graphical user interfaces [2] and conversation [18, 27].
While the utility of user-centred interactive explanations is evident in recent literature, the evaluation of these interactions remains a key research challenge. Current works primarily target the development of objective metrics for single-shot techniques [24, 28], emphasising the need for reproducible benchmarks on public datasets [16]. Such metrics are system-centred and model-agnostic, giving the advantage of generalisability. However, objective metrics fail to acknowledge the requirements of different stakeholder groups. A satisfactory explanation is reliant on the recipient’s expertise within that domain and of AI in general [20]. Subjective metrics, such as those described in [10, 11], allow an evaluation which is personalised to the individual and domain. However, existing subjective evaluations lack the capacity to measure the interactive process that underpins multi-shot explanations and how they impact user experience.
Table 1: Glossary
@@
We address this challenge by introducing the XAI Experience Quality (XEQ) Scale (pronounced: "seek" Scale). We define an XAI Experience as the user-centred process of a stakeholder interacting with an XAI system to gain knowledge and/or improve comprehension. XAI Experience Quality (XEQ) is defined as the extent to which a stakeholder’s explanation needs are satisfied by their XAI Experience. A glossary of all related terminology used throughout this paper is included in Table 1. Specifically, we ask the research question: “How to evaluate an XAI experience, in contrast to assessing single-shot (non-interactive) explanations?”. To address this, we follow a formal psychometric scale development process [3] and outline the following objectives:
@@
    perform a pilot study to refine and validate the XEQ scale for internal consistency, construct and discriminant validity.
The rest of this paper expands on each objective. We discuss related work in Section 2. Section 3 presents key previous publications and the creation of the initial items bank. The Content Validity study details and results are presented in Section 4 followed by Section 5 presenting pilot study details for the refinement and validation of the XEQ Scale. Key implications of the XEQ Scale are discussed in Section 6. Finally, we offer conclusions in Section 7.
@@ 2 Related Work
3 Literature Review and Initial Items Bank Compilation
@@ 3.1 Methodology
To scope the existing work and form the initial item bank, we conducted a targeted literature review in the domain of XAI evaluation metrics. The reasoning for a targeted review instead of a systematic review is two fold: 1) the purpose of the review is to form the initial item bank which involves in depth analysis of selected literature (depth over breadth); and 2) literature under this topic is significantly limited. The initial findings highlighted that while many publications discuss and emphasise the importance of evaluation dimensions (what should be or is evaluated), only a few actually propose and systematically develop metrics for XAI evaluation.
3.2 Findings: Evaluation Dimensions and Metrics
Hoffman et al. [9] are one of the leading contributors and their work has been widely utilised in many user-centred XAI research. They conceptually modelled the “process of explaining in XAI” outlining dimensions and metrics for evaluating single-shot explanations from stakeholders’ perspectives. They considered six evaluation dimensions: goodness, satisfaction, mental model, curiosity, trust and performance, For each dimension, they either systematically developed an evaluation metric or critiqued metrics available in literature offering a comprehensive evaluation methodology for XAI practitioners. System Causability Scale [11] is the other most prominent work in XAI evaluation. We discuss each scale briefly below.
@@ 3.2 Findings: Evaluation Dimensions and Metrics
was designed using psychometric theory to evaluate the subjective “goodness” of explanations with stakeholders. It consists of 8 items responded in a 5-step Likert Scale. It is viewed as the user-centred variant of the Goodness Checklist with many shared items. However, conversely, it has been evaluated for content validity with XAI experts as well as construct and discriminant validity in pilot studies.
@@ 3.2.1 Other Dimensions
Many other publications emphasised the need for user-centred XAI evaluations and explored evaluation dimensions. Two other dimensions considered in [9] are mental model and performance concerning task completion. Hoffman et al. recommended eliciting the mental model of stakeholders in think-aloud problem-solving and question-answering sessions. Performance is measured by observing the change in productivity and change in system usage. The evaluation of these dimensions requires metrics beyond questionnaire-based techniques. Another domain-agnostic survey finds many overlaps with Hoffman et al., defining 4 user-centred evaluation dimensions: mental model, usefulness and satisfaction, trust and reliance and human-task performance [20]. Zhou et al., [29] summarise previous literature, emphasising three subjective dimensions - trust, confidence and preference that overlap with dimensions identified in [9]. Conversely to Hoffman et al., they consider task completion to be an objective dimension in user-centred XAI evaluation.
@@
These findings highlighted that there are many overlaps between evaluation dimensions identified in recent years. However, we highlight two main gaps in this current work: 1) there is no consensus in previous literature regarding the applicable metrics to measure these evaluation dimensions; and 2) the majority of the existing dimensions and metrics focus on evaluating individual explanations, not the XAI experiences.
3.3 Compiling the Initial Item Bank
The initial item bank of 40 items included 7 from the Goodness Checklist [10]; 8 from the Satisfaction Scale [10]; 8 from the Trust Scale [10]; and 10 from the System Causability Scale [11]. Seven additional items were authored and included by the research team. These were designed to capture stakeholder views on the interactive experience which is less explicitly addressed in previous literature.
This initial list underwent a rigorous review and revision process, during which the research team eliminated duplicates, consolidated similar items, and rephrased items to reflect the measurement of XAI experiences instead of explanations. The resulting 32 statements formed the initial XEQ Scale (included in Supplementary Material). Response for each item is recorded on a 5-point Likert scale, ranging from “I Strongly Agree” to “I Strongly Disagree”.
@@ 3.3.1 Evaluation Dimensions
We reviewed evaluation dimensions from previous literature and consolidated XEQ items into four evaluation dimensions representing XAI experience quality: learning, utility, fulfillment, and engagement. These dimensions are relevant to capturing personalised experiences for a given stakeholder. We define them as follows:
@@ 4 XEQ Scale Development
This section presents the details and results of an expert user study performed to establish the content validity of the XEQ Scale.
4.1 Study Design
In this study, participants evaluated the initial set of 32 items using the Content Validity Ratio (CVR) method [15]. The CVR method is recommended for quantifying the strength of psychometric scale items with a small group of experts (5-10).
At the start, participants are familiarised with the terminology and they explore 3 sample XAI experiences: 1) a student interacting with a chatbot that assists with course recommendations and support; 2) a clinician interacting with the graphical interface of a radiograph fracture detection system for clinical decision support; and 3) a regulatory officer interacting with a local council welfare web page to explore the fairness and biases with the recommender system used for predicting application outcomes. A sample dialogue with the interactive Course Assist chatbot is shown in Figure 1. This dialogue, in its entirety, captures the explanation experience that we wish to evaluate. Next, participants are asked to rate the 32 items in terms of their relevance for measuring XAI experience quality using a 5-point Likert scale. The relevance scale ranges from Not Relevant at All to Extremely Relevant. Additionally, we included ratings related to clarity, which also use a Likert scale ranging from Not Clear at All to Extremely Clear. This clarity rating was added to get feedback on how easily the items are understood by the participants. Finally, we provided participants the opportunity to suggest rephrasing for items they found relevant but not clear.
4.2 Recruitment Details
The participants of this study consisted of XAI experts both from academia and the industry. 38 experts were contacted via email and 13 participated in the study.The 13 participants represented a diverse set of interests in the human-centred XAI research domain and are either actively conducting research or have published XAI research outcomes since 2020. The study was hosted on the Jisc Online Surveys platform for 3 weeks between November and December 2023.
4.3 Metrics
4.3.1 Content Validity
@@ 4.4 Results
S-CVI$(a)$ and S-CVI$(b)$ of the scale were 0.8846 and 0.2222. While S-CVI$(a)$ is comparable to the baseline of $0.9$, S-CVI$(b)$ indicate universal agreement is not achieved. However, existing literature suggests that meeting one of the baseline criteria is sufficient to proceed to pilot studies. Notably, the 14 items with I-CVI $\geq 0.78$ also only achieve average agreement (S-CVI$(a)=0.9179$) and not universal agreement (S-CVI$(b)=0.2667$). Following the item selection and refinement, each item was assigned an evaluation dimension based on the consensus of the research team (see Column “Factor” in Table 2). These will be used in further investigations using factor analysis to establish the construct validity of the scale.
@@ 5 XEQ Scale Refinement and Validation
In this section, we present the pilot study conducted to refine the XEQ Scale for internal consistency, construct validity and discriminant validity.
5.1 Study Design and Applications
The study involved two application domains: 1) CourseAssist Chatbot for new students to guide their course selection processes and 2) AssistHub website for welfare applicants to assist with application outcomes. For each application, two sample XAI experiences were designed one relatively positive and one relatively negative experience. The two samples differ in the quality of the explanations presented to the participant and the resulting impact on the overall interaction flow. Participants accessed all samples in video format. Figure 1 presents the static views of the two samples of the CourseAssist Chatbot. AssistHub samples are included in Supplementary Material. These sample experiences create a controlled experiment where the discriminant properties of the scale can be validated.
First, the participants explore the XAI experience and they proceed to respond to the XEQ Scale. In addition, they are also queried about the clarity of items within the scope of the sample experience. Lastly, participants can offer free-text feedback about the study.

### v4

@@ Abstract
Explainable Artificial Intelligence (XAI) aims to improve the transparency of autonomous decision-making through explanations. Recent literature has emphasised users’ need for holistic “multi-shot” explanations and personalised engagement with XAI systems. We refer to this user-centred interaction as an XAI Experience. Despite advances in creating XAI experiences, evaluating them in a user-centred manner has remained challenging. In response, we developed the XAI Experience Quality (XEQ) Scale. XEQ quantifies the quality of experiences across four dimensions: learning, utility, fulfilment and engagement. These contributions extend the state-of-the-art of XAI evaluation, moving beyond the one-dimensional metrics frequently developed to assess single-shot explanations. This paper presents the XEQ scale development and validation process, including content validation with XAI experts, and discriminant and construct validation through a large-scale pilot study. Our pilot study results offer strong evidence that establishes the XEQ Scale as a comprehensive framework for evaluating user-centred XAI experiences.
keywords:
@@ 1 Introduction
Explainable Artificial Intelligence (XAI) describes a range of techniques to elucidate autonomous decision-making and the data that informed that AI system [1, 2, 3]. Each technique typically provides explanations focusing on a specific aspect of the system and its decisions, often answering a singular question fulfilling a specific intent [4, 5]. Accordingly, the utility of employing multiple techniques for a holistic explanation of a system becomes increasingly clear [3, 6]. The collection of explanations, provided by different techniques and describing different components of the system, forms what we describe as “multi-shot” explanations. Multi-shot explanation treats the elucidation of an AI system as a process which can invoke a variety of knowledge sources to construct explanations that comprehensively answer a user’s explanation needs. This is unlike single-shot, which provides a lone explanation that targets a specific aspect of an autonomous decision to resolve an individual query (see Figure 1). Furthermore, viewing multi-shot explanations as an interactive process allows users to engage through graphical interfaces [3] or conversations [7, 6], enabling personalised experiences tailored to individual needs [8].
Figure 1: A conversational interaction demonstrating a single-shot explanation (adapted from IBM Explainability 360 platform [3]) extended to a multi-shot XAI experience. Each single-shot explanation is represented by individual turns in the conversation, while a multi-shot explanation is represented by the conversation in its entirety.
While the utility of user-centred interactive multi-shot explanations is evident in recent literature, evaluating them remains a key research challenge. Current works primarily target the development of objective metrics for single-shot techniques [9, 10], emphasising the need for reproducible benchmarks on public datasets [11]. Such metrics are system-centred and model-agnostic, meaning they are compatible with a variety of XAI systems and giving the advantage of generalisability. However, objective metrics fail to acknowledge the requirements of different stakeholder groups. A satisfactory explanation is reliant on the recipient’s expertise within that domain [12] and their previous experiences with AI [13, 14]. Subjective metrics, such as those described in [15, 16], allow an evaluation which is personalised to the individual and domain. However, existing subjective evaluations lack the capacity to measure the interactive process that underpins multi-shot explanations and how they impact user experience. There is a clear need for a tool that facilitates the evaluation of XAI experiences. In turn, this would empower the iterative development of XAI systems to ensure they address the diverse needs of various stakeholder groups.
Table 1: Glossary of term and definitions used throughout this paper.
@@
Single-Shot Explanation An explanatory process where a lone explanation is provided to answer a specific query.
Multi-Shot Explanation An explanatory process where multiple techniques are used to provide two or more alternative and/or complementary insights into a single XAI system or output.
@@
We address this challenge by introducing the XAI Experience Quality (XEQ) Scale. We define an XAI Experience as the user-centred process of a stakeholder interacting with an XAI system to gain knowledge and/or improve comprehension. XAI Experience Quality (XEQ) is defined as the extent to which a stakeholder’s explanation needs are satisfied by their XAI Experience. A glossary of all related terminology used throughout this paper is included in Table 1. Specifically, we ask the research question: “How to evaluate an XAI experience, in contrast to assessing single-shot (non-interactive) explanations?”. To address this, we follow a formal psychometric scale development process [17] and outline the following objectives:
@@
    perform a pilot study to refine and validate the XEQ scale for internal consistency, discriminant and construct validity and test-retest reliability.
The rest of this paper expands on each objective. We discuss related work in Section 2. Section 3 presents the creation of the initial items bank. The Content Validity study details and results are presented in Section 4 followed by Section 5 presenting the pilot study for the refinement and validation of the XEQ Scale. Key implications and the practical use of the Scale are discussed in Section 6. Finally, we offer conclusions in Section 7.
@@ 2 Related Work
3 Initial Items Bank Compilation
@@ 3.1 Methodology
Netemeyer [23] outlines a framework for developing an item pool, via domain sampling grounded literature review. Following this approach, we conducted a targeted literature review to examine existing research, identify key dimensions, and develop the initial item bank. The reasoning for a targeted review instead of a systematic review is twofold: 1) the purpose of the review is to form the initial item bank which involves in-depth analysis of selected literature (depth over breadth); and 2) literature under this topic is significantly limited. The initial findings highlighted that while many publications discuss and emphasise the importance of evaluation dimensions (what should be or is evaluated), only a few propose and systematically develop metrics for XAI evaluation.
3.2 Findings: Evaluation metrics and dimensions
Hoffman et al. [15] are one of the leading contributors and their work has been widely utilised in many user-centred XAI research. They conceptually modelled the “process of explaining in XAI” outlining dimensions and metrics for evaluating single-shot explanations from stakeholders’ perspectives. They considered six evaluation dimensions: goodness, satisfaction, mental model, curiosity, trust and performance, For each dimension, they either systematically developed an evaluation metric or critiqued metrics available in literature offering a comprehensive evaluation methodology for XAI practitioners. System Causability Scale [16] is the other most prominent work in XAI evaluation. Notably, all the above scales are designed to be application-domain agnostic, with each item adaptable to reference the specific AI system under evaluation. We discuss each scale briefly below.
@@ 3.2 Findings: Evaluation metrics and dimensions
was designed using psychometric theory to evaluate the subjective “goodness” of explanations with target users. It consists of 8 items responded in a 5-step Likert Scale. It is viewed as the user-centred variant of the Goodness Checklist with many shared items. This scale has been evaluated for content validity with XAI experts as well as construct and discriminant validity in pilot studies.
@@ 3.2.1 Other Dimensions
Many other publications emphasised the need for user-centred XAI evaluations and explored evaluation dimensions. Two other dimensions considered by Hoffman et al. [15] are mental model and performance concerning task completion. They recommended eliciting the mental model of stakeholders in think-aloud problem-solving and question-answering sessions. Performance is measured by observing the change in productivity and change in system usage. The evaluation of these dimensions requires metrics beyond questionnaire-based techniques. Another domain-agnostic survey finds many overlaps with Hoffman et al., defining 4 user-centred evaluation dimensions: mental model, usefulness and satisfaction, trust and reliance and human-task performance [12]. Zhou et al., [26] summarise previous literature, emphasising three subjective dimensions - trust, confidence and preference that overlap with dimensions identified in [15]. Conversely to Hoffman et al., they consider task completion to be an objective dimension in user-centred XAI evaluation.
@@
These findings highlighted that there are many overlaps between evaluation dimensions identified in recent years. However, we highlight two main gaps in this current work: 1) there is no consensus in previous literature regarding the applicable metrics to measure these evaluation dimensions; and 2) the majority of the existing dimensions and metrics focus on evaluating individual explanations (i.e. single-shot), not multi-shot XAI experiences.
3.3 Initial item bank
Informed item sampling from literature resulted in 33 items, including 7 from the Goodness Checklist [15]; 8 from the Satisfaction Scale [15]; 8 from the Trust Scale [15]; and 10 from the System Causability Scale [16]. As evidenced in the literature review, while these items covered constructs related to user-centred evaluation of single-shot explanations, they lacked explicit representativeness of multi-shot explanations. Recognising that interaction is crucial to creating effective explanation experiences, the research team generated seven additional items to capture this construct. These items were designed to capture stakeholder perspectives on the interactive experience, a dimension explicitly not covered in previous XAI evaluation literature. This expanded the item pool to a total of 40 items.
The item pool was then standardised to a symmetric 5-point Likert Scale, ranging from “I Strongly Agree” to “I Strongly Disagree”. The existing user-centred XAI scales identified in the literature review commonly adopt the 5-point format, offering a balance between granularity and minimising ambiguity between response options. Broader research in psychology has also shown via mathematical modelling that using a scale with more than five points offers no significant improvement in internal consistency reliability measured by Cronbach’s alpha [29].
Finally, following similar approaches from recent literature in scale development [30, 31], the item pool then underwent a 2-stage rigorous review and revision process. Each stage of the revision process was led by two authors, who proposed revisions, and subsequently validated by the remaining authors, who either accepted, rejected, or further revised them. In the first stage, lexically identical items were removed, and items with significant semantic overlap were consolidated, resulting in 32 items. During the second stage, expressions were made more concise, non-suggestive, and aligned with the intended construct (i.e., measuring XAI experiences rather than explanations). Rephrased items were further validated by iterating wording until unanimous agreement between authors to ensure clarity for stakeholders with varying knowledge levels and to eliminate suggestive wording to minimise potential bias. The resulting 32 items formed the initial XEQ Scale (included in Supplementary Material).
@@ 3.3.1 Evaluation Dimensions
We reviewed evaluation dimensions from previous literature and consolidated XEQ items into four evaluation dimensions representing XAI experience quality: learning, utility, fulfilment, and engagement. These dimensions are relevant to capturing personalised experiences for a given stakeholder. We define them as follows:
@@ 4 XEQ Scale Development
This section presents the details and results of the expert-led evaluation of the initial set of 32 items to develop the XEQ scale, following the Content Validity Ratio (CVR) method [33, 34].
4.1 Study instrument
The study instrument consisted of three components: 1) an introduction to terminology; 2) three XAI experience examples; and 3) a structured survey for data collection. The first component introduced the participants to the terminology used throughout the study (also included in Table 1). This is intended to establish a consistent mental model between experts from diverse XAI backgrounds. Next, participants explored three XAI experiences curated to capture different application domains, end-user explanation needs and interaction modalities. Lastly, participants responded to the survey with three sub-sections: a) rating the 32 items regarding their relevance for measuring XAI experience quality on a 5-step Likert scale ranging from Not Relevant at All to Extremely Relevant; b) rating the 32 items regarding their clarity on a similar Likert scale ranging from Not Clear at All to Extremely Clear; and c) rephrasing items that they found relevant but not clear.
4.1.1 XAI experiences development and selection
The development of sample XAI experiences was conducted over three stages. First, we established several desiderata for sample experiences: a) reflect diverse, real-world AI systems where explaining AI outcomes is critical; b) possess the level of complexity that is suited for multi-shot explanation; and c) capture varying levels of domain expertise and AI familiarity.
The second stage builds on these desiderata, to identify relevant application domains for sample XAI experiences considering previous literature on XAI design and development, and feedback from industry and academic partners. We selected radiograph fracture detection from the medical domain, welfare application approval auditing from the governance domain, and student support for course selection from academia. Regulations governing the use of AI decision-support in medical imaging underscore the need for XAI, particularly through comprehensive explanation methodologies [35, 36, 37, 38]. Similar regulations are being implemented for AI in civic and governmental decision-making [39] to ensure that automated decisions are interpretable, contestable, auditable, and fair [40]. Finally, student support represents an area with significantly lower legal/ethical requirements for explainability, but with high user experience impact [41].
In the third and final stage, we designed each of the sample experiences. For each experience, the target user was identified to ensure a variety of AI knowledge and domain expertise levels. The interaction modality was selected to suit the target user’s explanation and interaction needs. The sample interactions articulated different types of explanations addressing the target user’s explanation needs. The three sample experiences are summarised in Table 2. A sample dialogue with the interactive CourseAssist chatbot is shown in Figure 2 positive XAI experience. This dialogue, in its entirety, captures an XAI experience that we wish to evaluate. The experiences constructed for medical imaging and welfare application approval auditing are available in the Supplementary Material.
Table 2: Sample XAI experiences; for each experience, we summarise the AI system, attributes of the target stakeholder and the explanations presented in their XAI experience. Knowledge levels are adapted from the Dreyfus 6-stage skill acquisition model [42].
Identifier Attributes Name Domain knowledge AI familiarity Interaction Modality
CourseAssist Stakeholder Student Novice Novice Conversation
AI System A rule-based agent that recommends courses given a student’s chosen career path
Explanations Distribution and statistics of the data used in the recommendations;
Factual explanations; Nearest-neighbour cases
RadioAssist Stakeholder Trainee Radiologist Proficient Competent Graphical User Interface
AI System A neural network black-box that predicts the presence of fracture given
a radiograph in PNG format
Explanations Distribution and statistics of the data used to develop the model; Feature attributions;
Nearest-neighbour cases; Performance metrics of the AI model
AssistHub Stakeholder Auditor Expert Proficient Graphical User Interface
AI System A decision tree algorithm that recommends the outcome given a welfare application
Explanations Distribution and statistics of the data used to create the AI system;
Prototypical approved and rejected cases; Outlier cases; Simplified decision tree
4.2 Participant recruitment and selection
The CVR method recommends a small group of domain experts (5-10) for quantifying the strength of psychometric scale items [33]. We invited 38 XAI experts with diverse expertise within XAI, considering those who are actively publishing in the XAI domain since 2020. The study recruited a diverse group of experts, 3 from industry and 10 from academia. Their specific areas of expertise were identified as follows: 3 experts in implementing XAI in industrial engineering applications (Energy and Telecommunications); 1 expert in XAI evaluation scale development; 5 experts in example-based and counterfactual XAI methods; 2 experts in human-centric XAI design and development; and 2 experts in the role of causality in XAI. They collectively represent the diverse aspects of crafting effective XAI experiences.
4.3 Data collection and analysis
The study was hosted on the Jisc Online Surveys platform for 3 weeks and collected responses from 13 XAI experts. Their responses were collated and analysed using the following metrics.
Content Validity
@@ 4.4 Results
S-CVI$(a)$ and S-CVI$(b)$ of the scale were 0.8846 and 0.2222. While S-CVI$(a)$ is comparable to the baseline of $0.9$, S-CVI$(b)$ indicate universal agreement is not achieved. However, existing literature suggests that meeting one of the baseline criteria is sufficient to proceed to pilot studies. Notably, the 14 items with I-CVI $\geq 0.78$ also only achieve average agreement (S-CVI$(a)=0.9179$) and not universal agreement (S-CVI$(b)=0.2667$).
Following the item selection, each item was assigned an evaluation dimension based on the consensus of the research team (see Column “Evaluation Dimension” in Table 4). These will be used in further investigations using confirmatory factor analysis in Section 5 to establish the construct validity.
@@ 5 XEQ Scale Refinement and Validation
In this section, we present the pilot study conducted to refine the XEQ Scale for internal consistency and construct validity, discriminant validity and test-retest reliability.
5.1 Study instrument
The pilot study instrument consisted of three components: 1) an introduction to the terminology; 2) the explanation experience; and 3) a structured survey for data collection. Similar to the CVR study, introducing the terminology is intended to establish a common understanding among participants. Next participants explored a sample XAI experience stepping into the role of the target user. Finally, participants evaluated the sample XAI experience by responding to the 18 items each on a 5-step Likert scale ranging from I strongly agree to I strongly disagree.
5.1.1 XAI experiences development and selection
The pilot study is large-scale (100+ participants) and aims, among other goals, to establish discriminant validity. To achieve this, we identified the following practical considerations for designing sample experiences: a) ensuring access to a large pool of target users via survey recruitment platforms or within the research institute; b) curating a relatively negative experience counterpart for each sample for validating the scale’s discriminant properties in a controlled setting; and c) ensuring access to the same sample over longitudinal follow-up to validate the scale’s test-retest reliability.
Based on these considerations, we selected the CourseAssist and AssistHub systems for sample experiences from the point of view of a student and a welfare applicant respectively. Negative experiences, in contrast to positive ones, were deliberately designed to lack best practices recommended in recent human-centred XAI literature [43, 44, 7, 45]. Specifically, they lacked: a) diversity in explanations within the experience; b) alignment between the target user’s intent and the explanations provided; c) alignment between the user’s knowledge level and the explanations; and d) mechanisms to encourage engagement with the XAI agent. The resulting four sample experiences are identified as CourseAssist-P, CourseAssist-N, AssistHub-P, AssistHub-N respectively where P and N indicate positive and negative experiences. Figure 2 presents the static views of CourseAssist-P and CourseAssist-N. AssistHub samples are included in Supplementary Material.